OpenStudy (zasharra):

This won't take long! Need help fast. (May 14, 1999) Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (zasharra):

@ganeshie8 @Hero @pooja195 @Michele_Laino @robtobey @ash2326

OpenStudy (zasharra):

@AravindG @uri @mathmale @misty1212 @KendrickLamar2014 @bibby

OpenStudy (zasharra):

@karolynxx @kai47 @AskMeLotss @kimbllse777 @Vickyisthesmartone123

OpenStudy (zasharra):

Guys?

OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. f(x)=3(x-8)2+15 A. g(x)=2(x+-8)2+15. This meets the requirements because it opens in the same direction as function f(x) because a is positive which means it opens upward. It has the same vertex, 10, 1 because the h and k are the same as the function f(x), and it will be narrower since a is 2 which is less than the 3 used in the function f(x). B. h(x)=-2(x-8)2+15. This meets the requirements because it will open in the opposite direction, downwards, since a is negative. It has the same vertex, 10, 1 because the h and k are the same as the function f(x), and it will be wider because -2 is less than 3 and the smaller the number, the wider the parabola.

OpenStudy (anonymous):

Google how to change an equation from vertex to standard form for #2

OpenStudy (anonymous):

For #5 plug in 4 and 7 and solve like you would for a function and then find the slope for them which would give you the rate of change

OpenStudy (anonymous):

If you copy and paste those functions into this website: http://www.wolframalpha.com/ It will give you the parabola and you can save the image and then post it into your ducument

OpenStudy (zasharra):

@AskMeLotss my birthday 05/14/99

OpenStudy (anonymous):

I dont get it...

OpenStudy (zasharra):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below.

OpenStudy (anonymous):

ya

OpenStudy (zasharra):

nvm i guess

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

f(x)=3(x-8)2+15

OpenStudy (anonymous):

hope I helped you, I am trying to get help on my math problems...its hard and I hate it :(

OpenStudy (zasharra):

yeah i understand thanks anyways

OpenStudy (anonymous):

your welcome...XD